# Real-Time Machine Learning Strategies for a New Kind of Neuroscience Experiments

### v1

@@ I Introduction
This article is accepted for publication in the 2024 European Signal Processing Conference (EUSIPCO)
Offline analysis of data recorded during an experiment has become the standard paradigm in neuroscience for investigating the neural underpinnings of behavior. These analyses often reveal the need to revise experimental conditions, or deficiencies in the data collection process itself, potentially rendering it uninformative about the scientific hypothesis under consideration. Consequently, experimentalists carry out multiple iterations of an experiment to obtain definitive findings, making this mode of analysis expensive and time-consuming. The scientific questions that can be addressed from pre-collected data are also limited, especially for establishing causal links between neural activity and behavior. With advances in recording techniques over the past few decades, we have access to a large population of simultaneously recorded neurons during an experiment. The shift from studying single neurons to population level activity has allowed us to probe more intricate neural phenomena and has significantly enhanced our understanding of neural computation. However, the increased spatiotemporal complexity of these recordings has made real-time analysis a challenging endeavor.
@@ III-A Convergence and Sample Efficiency
As discussed in Section II, SSMs can be readily used to infer the latent state and learn the parameters of the generative model from streaming data. However, online estimation necessitates constant space and time complexity and the optimization of these models is often non-convex and data intensive. This is because unlike offline training, each recorded sample is processed only once during optimization and consequently, a substantial number of samples are required for these models to attain operational efficacy. This results in undesirable experimental costs and hampers practical application, as experimental durations are typically restricted. While these issues could be partly alleviated by incorporating prior knowledge in the form of parameter initializations or optimal hyperparameter values, these models are typically trained from scratch due to a lack of inductive biases for neural dynamics.
@@ IV-A Large-scale Integrative Neuroscience
Apart from improving scientific understanding, this integrated model can be useful for constraining the search space of online filtering algorithms, thereby reducing the problem complexity. For instance, it is widely known that the neural dynamics underlying neural activity in motor cortex across animals engaged in similar tasks show qualitative similarities as well [32]. This has been leveraged to transfer knowledge in the form of expected latent dynamics across recordings [33]. Large-scale integrative models can provide further inductive biases for diverse settings and enable faster convergence of online filtering algorithms. This inductive biases can complement biologically informed constraints and parameterizations of latent dynamics to improve identifiability of learnt solutions [34, 35].
@@ IV-B Meta Learning
Meta-learning approaches have been widely used in machine learning to train models that are capable of generalizing to novel settings with few training samples [36, 37]. However, there has been limited work on applying these methods for improving the convergence speed and sample efficiency of online filtering methods.
There are two broad categories of meta-learning approaches – gradient-based methods such as model agnostic meta-learning learn parameter initializations for faster convergence on downstream tasks [38]; in contrast, model-based approaches train a black-box meta-learner along with the base model to directly generate parameters of the base model [39, 40]. These methods can be applied to meta-learn parameter initializations, optimal learning rate, latent state dimensionality, etc., by optimizing the performance of the SSM on diverse, pre-collected recordings offline. This trained meta-learner can subsequently be fine-tuned at the beginning of the experiment to provide data-driven priors to online filtering algorithms (Fig. 1). While this offers a promising solution for improving model convergence and sample efficiency, existing meta-learning approaches need to be adapted to deal with statistical heterogeneities across different neural datasets, for instance, differences in the number and tuning properties of recorded neurons, the dimensionality of neural state, and so on.
@@ IV-C Novel Control Methods
Apart from enabling system identification, the development of control objectives is imperative for investigating complex scientific phenomena. Conventionally, model predictive control (MPC) employs the current state, observations and external inputs to predict the system’s next state. A longer prediction horizon is achieved by recursively feeding the model’s output back as the new input, along with the remaining variables, in a rolling horizon forecast fashion. This approach, while straightforward in theory, is both costly, as it requires the entire model to perform multiple sequential inferences, and susceptible to accruing inaccuracies over time, as it will propagate its own prediction error. We foresee an opportunity in MPC to leverage the advances in differentiable and computationally efficient sequence-to-sequence modeling that revolutionized language models. For example, causal transformers could infer latent processes[43] and serve as inverse controllers in an autoencoder framework (Fig. 3). The inference and control problems can be parallelizable and the latency can be easily reduced via scaling computing hardware, enabling real-time control sequences generation for neural perturbation.

### v2

@@ I Introduction
The article includes minor revisions of the paper presented at the 2024 European Signal Processing Conference (EUSIPCO), Lyon, France.
Offline analysis of data recorded during an experiment has become the standard paradigm in neuroscience for investigating the neural underpinnings of behavior. These analyses often reveal the need to revise experimental conditions, or deficiencies in the data collection process itself, potentially rendering it uninformative about the scientific hypothesis under consideration [1]. Consequently, experimentalists carry out multiple iterations of an experiment to obtain definitive findings, making this mode of analysis expensive and time-consuming. The scientific questions that can be addressed from pre-collected data are also limited, especially for establishing causal links between neural activity and behavior [2]. With advances in recording techniques over the past few decades, we have access to a large population of simultaneously recorded neurons during an experiment. The shift from studying single neurons to population level activity has allowed us to probe more intricate neural phenomena and has significantly enhanced our understanding of neural computation. However, the increased spatiotemporal complexity of these recordings has made real-time analysis a challenging endeavor.
@@ III-A Convergence and Sample Efficiency
As discussed in Section II, SSMs can be readily used to infer the latent state and learn the parameters of the generative model from streaming data. However, online estimation necessitates constant space and time complexity and the optimization of these models is often non-convex and data intensive [22, 13]. This is because unlike offline training, each recorded sample is processed only once during optimization and consequently, a substantial number of samples are required for these models to attain operational efficacy. This results in undesirable experimental costs and hampers practical application, as experimental durations are typically restricted. While these issues could be partly alleviated by incorporating prior knowledge in the form of parameter initializations or optimal hyperparameter values, these models are typically trained from scratch due to a lack of inductive biases for neural dynamics.
@@ IV-A Large-scale Integrative Neuroscience
Apart from improving scientific understanding, this integrated model can be useful for constraining the search space of online filtering algorithms, thereby reducing the problem complexity. For instance, it is widely known that the neural dynamics underlying neural activity in motor cortex across animals engaged in similar tasks show qualitative similarities as well [34]. This has been leveraged to transfer knowledge in the form of expected latent dynamics across recordings [35]. Large-scale integrative models can provide further inductive biases for diverse settings and enable faster convergence of online filtering algorithms. These inductive biases can complement biologically informed constraints and parameterizations of latent dynamics to improve identifiability of learnt solutions [36, 37].
@@ IV-B Meta Learning
Meta-learning approaches in machine learning aim to train models that generalize to novel settings with few training samples [38, 39]. However, there has been limited work on applying these methods for improving the convergence speed and sample efficiency of online filtering methods.
There are two broad categories of meta-learning approaches – gradient-based methods such as model agnostic meta-learning learn parameter initializations for faster convergence on downstream tasks [40]; in contrast, model-based approaches train a black-box meta-learner along with the base model to directly generate parameters of the base model [41, 42]. These methods can be applied to meta-learn parameter initializations, optimal learning rate, latent state dimensionality, etc., by optimizing the performance of the SSM on diverse, pre-collected recordings offline. This trained meta-learner can subsequently be fine-tuned at the beginning of the experiment to provide data-driven priors to online filtering algorithms (Fig. 1). While this offers a promising solution for improving model convergence and sample efficiency, existing meta-learning approaches need to be adapted to deal with statistical heterogeneities across different neural datasets, for instance, differences in the number and tuning properties of recorded neurons, the dimensionality of neural state, and to incorporate prior knowledge and structural hypotheses.
@@ IV-C Novel Control Methods
In addition to enabling system identification, the development of control objectives is imperative for investigating complex scientific phenomena. Conventionally, model predictive control (MPC) employs the current state, observations and external inputs to predict the system’s next state. A longer prediction horizon is achieved by recursively feeding the model’s output back as the new input, along with the remaining variables, in a rolling horizon forecast fashion. This approach, while straightforward in theory, is both costly, as it requires the entire model to perform multiple sequential inferences, and susceptible to accruing inaccuracies over time, as it will propagate its own prediction error. We foresee an opportunity in MPC to leverage the advances in differentiable and computationally efficient sequence-to-sequence modeling that revolutionized language models. For example, causal transformers could infer latent processes[45] and serve as inverse controllers in an autoencoder framework (Fig. 3). The inference and control problems can be parallelized and the latency can be reduced via scaling computing hardware, enabling better real-time control through faster generation of candidate sequences for neural perturbation.